If the line joining circumcentr and incentre is parallel to BC then cosB+cosC=

Grade:12

4 Answers

P srivadhan
8 years ago
SO ANSWER IS I.E COSB+COSC=1 HERE this angle theta is the angle between IO and LO it is zero in this case so our answer is one.
P srivadhan
8 years ago
I had seen that above image submitted by me is unclear . so to view the solution please click the below link . I don’t see ny option to attach image. copy it and see in broswer. askiitans1.weebly.com/link-1.html
P srivadhan
7 years ago
askiitans1.weebly.com/link-1.html.
 
 
 
I had seen that above image submitted by me is unclear . so to view the solution please click the below link . I don’t see ny option to attach image. copy it and see in broswer. askiitans1.weebly.com/link-1.html
mycroft holmes
7 years ago
Drop perpendiculars from I and O on to BC and let the feet of the perpendiculars be P and Q resp.
 
IP = in-radius = r.
 
Also OQ = R cos A.
 
Since OI is parallel to BC, we have IP = OQ i.e. r = R cos A and hence
 
 1+\frac{r}{R} = 1+\cos A
 
We know that 1+\frac{r}{R} = \cos A + \cos B + \cos C
 
Thus we obtain 1+\cos A = \cos A + \cos B + \cos C \Rightarrow \boxed {\cos B+\cos C =1}